# NoMIRACL: Knowing When You Don’t Know for Robust Multilingual Retrieval-Augmented Generation

### 18 Dec 2023 v1

@@ Abstract
Retrieval-augmented generation (RAG) grounds large language model (LLM) output by leveraging external knowledge sources to reduce factual hallucinations. However, prior works lack a comprehensive evaluation of different language families, making it challenging to evaluate LLM robustness against errors in external retrieved knowledge. To overcome this, we establish NoMIRACL, a human-annotated dataset for evaluating LLM robustness in RAG across 18 typologically diverse languages. NoMIRACL includes both a non-relevant and a relevant subset. Queries in the non-relevant subset contain passages manually judged as non-relevant or noisy, whereas queries in the relevant subset include at least a single judged relevant passage. We measure LLM robustness using two metrics: (i) hallucination rate, measuring model tendency to hallucinate an answer, when the answer is not present in passages in the non-relevant subset, and (ii) error rate, measuring model inaccuracy to recognize relevant passages in the relevant subset. We build a GPT-4 baseline which achieves a 33.2% hallucination rate on the non-relevant and a 14.9% error rate on the relevant subset on average. Our evaluation reveals that GPT-4 hallucinates frequently in high-resource languages, such as French or English. This work highlights an important avenue for future research to improve LLM robustness to learn how to better reject non-relevant information in RAG. [^1] NoMIRACL dataset and the GPT-4 baseline are available at: https://github.com/project-miracl/nomiracl
@@ 1 Introduction
Retrieval Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021; Borgeaud et al., 2022) is a promising way to incorporate external knowledge via a first-stage retrieval system. RAG instills information from reliable knowledge corpora (provided as external passages) to generate accurate and faithful responses Shuster et al. (2021); Gao et al. (2023b); Liu et al. (2023). Ever since the introduction of large language models (LLMs), such as GPT-3 Brown et al. (2020) or LLAMA-2 Touvron et al. (2023), they are the de-facto choice for generation in RAG, due to their impressive performance across multiple reasoning and natural language inference tasks Chang et al. (2023); Guo et al. (2023). RAG grounds the LLM-generated output, thereby avoiding previously observed factual hallucination Maynez et al. (2020); Raunak et al. (2021) and outdated knowledge Cao et al. (2021); He et al. (2023) in LLMs.
@@
Figure 1: LLM robustness evaluation as a binary tree in NoMIRACL. When dealing with queries in the non-relevant subset ($\mathrm{F}$), the LLM is expected to disregard all noisy passages and refrain from answering ($\mathrm{N}$). Conversely, for queries in the relevant subset ($\mathrm{T}$), the LLM should recognize the relevant passage and provide a valid answer ($\mathrm{P}$).
@@
A challenging issue in RAG is to provide robust and reliable LLM-generated output. The generation stage is dependent on the first-stage information retrieval system. The retrieval system poses challenges in accurately retrieving relevant information when evaluated either on zero-shot domains Thakur et al. (2021) or low-resource languages Zhang et al. (2023). The incorrect or non-relevant information contained in retrieved passages can mislead the LLM to hallucinate Adlakha et al. (2023); Chen et al. (2023); Shi et al. (2023); Yoran et al. (2023); Yu et al. (2023). However, prior work lacks a comprehensive evaluation of LLM reasoning capabilities in multiple languages. As a result, it is unknown whether the LLM and its utilization of the external information is robust against first-stage retrieval errors across both high- or low-resource languages.
@@
In this work, we introduce NoMIRACL, a new dataset to evaluate LLM robustness against errors in first-stage external information, i.e., retrieved passages, across 18 typologically diverse languages. We hire over 30 native speakers as annotators Zhang et al. (2023) to construct the dataset. NoMIRACL contains two subsets, non-relevant and relevant. Where, the non-relevant subset contains queries with no-known answers, i.e., all top-$k$ retrieved passages manually judged to non-relevant. Conversely, the relevant subset contains queries with known answers, i.e., at least one of the top-$k$ passages is manually judged relevant.
We build an initial GPT-4 baseline Brown et al. (2020), a closed-book multilingual LLM on NoMIRACL using the top-$k$ oracle passages retrieved using a hybrid retrieval system from a language-specific Wikipedia corpus Zhang et al. (2023). We use a vanilla prompt template for GPT-4. Our results show that GPT-4 frequently tends to confuse non-relevant passages as relevant and achieves a high 33.2% hallucination rate on average on the non-relevant subset. On the other hand, GPT-4 achieves a lower 14.2% error rate on average on the relevant subset indicating a strong ability to identify relevant passages with known answers.
Furthermore, we observe a positive correlation (Spearman $\rho$ = 0.39) between the GPT-4 hallucination rate and the language resource size, i.e., Wikipedia corpus size Zhang et al. (2023). We suspect the information required to answer a query is better available in a high-resource language (i.e., harder negative), which potentially confuses the LLM to judge the passage as relevant. These results highlight the need to explore better techniques in future to improve LLM robustness.
@@
Figure 2: Contingency table for robustness evaluation with NoMIRACL. More details are provided in (§2.1); (Subset) denotes the ground-truth in NoMIRACL; (Pred.) denotes the LLM output prediction.
@@
To summarize our contributions are: (i) We introduce NoMIRACL, a novel multilingual dataset to assess LLM robustness against first-stage retrieval errors. (ii) We provide GPT-4 as a baseline and observe challenges in over-reliance on noisy passages in the non-relevant subset over recognizing relevant passages in the relevant subset; We hope NoMIRACL can serve as a valuable dataset towards a much-needed LLM robustness evaluation across diverse languages in RAG.
@@ 2 Background and Problem Identification
A challenging issue in RAG is to provide robust and reliable LLM-generated output against first-stage information retrieval system. Overreliance of the LLM on the content of the retrieved passages (i.e., tendency to extract information from passages) can be limiting when passages are noisy or non-relevant Yu et al. (2023); Yoran et al. (2023); Shi et al. (2023).
@@ 2.1 Robustness Evaluation
We conduct our robustness evaluation as a contingency table to robustly evaluate the LLM behavior in different scenarios. We evaluate a binary classification task, by comparing LLM predictions against the ground truth provided in NoMIRACL. We provide our contingency table in Figure 2.
@@
Confusion Matrix. In our confusion matrix (cf. Figure 2), False Negative ($\mathrm{FN}$) denotes when the model correctly predicts queries with no-known answers using the noisy retrieved passages, whereas False Positive ($\mathrm{FP}$) denotes when the model prediction is incorrect on the non-relevant subset. Similarly, True Positive (TP) denotes when LLM correctly predicts queries with known answers within the retrieved passages, whereas False Positive (FP) denotes when the model prediction is incorrect (model finds no answer) on the relevant subset.
Metrics. Following prior works Adlakha et al. (2023); Chen et al. (2023), we assess LLM robustness as a binary classification task using two metrics: (i) hallucination rate and (ii) error rate. First, we compute the hallucination rate (in %) $=\mathrm{FP}/\mathrm{FP}+\mathrm{FN}$ which measures the model tendency to hallucinate an answer, when no answer is available in all of the passages in the non-relevant subset. Next, we measure the error rate (in %) $=\mathrm{TN}/\mathrm{TN}+\mathrm{TP}$ which measures the model ability to identify the relevant passages in the relevant subset.
@@ 3 NoMIRACL Dataset
In this section, we describe our NoMIRACL construction procedure in (§3.1), fold creation in (§3.2), difficulty in providing answers for queries in NoMIRACL in (§3.3), and languages covered and dataset usage in (§3.4). An overview of our data construction procedure can be found in Figure 3.
@@ 3.1 Data Construction Procedure
NoMIRACL, a supplementary dataset, was created in the course of developing MIRACL Zhang et al. (2023). The data annotation procedure is conducted in two stages, following Zhang et al. (2023); Clark et al. (2020). In the first stage, the annotator (a native language speaker) is asked to generate a well-formed query for each individual prompt text. Each prompt is a short text snippet containing the first 100 words from a language-specific Wikipedia corpus $C$. Next, for each human-generated query, top-$k$ passages are retrieved from the Wikipedia corpus $C$ using a hybrid multilingual retrieval system (more details in §4.1). In the next stage, annotators assess the relevance of the top-$k$ query–passage pairs as a binary judgment, either relevant (relevance = $1$) or non-relevant (relevance = 00). For more details regarding the annotation procedure, such as the hiring process or stipend, we would like to refer the reader to Zhang et al. (2023).
Non-relevant subset. Annotators generate queries based on certain dataset guidelines, however, occasionally the generated queries cannot be answered with external corpus, which leads to the scenario where none of the top-$k$ passages are relevant, i.e., contains the answer. These queries with no-known answers occur during an annotation procedure due to the following reasons: (i) Queries can be either generic or specific for information to be present in Wikipedia, for e.g. “What does the AC button on a calculator stand for?” retrieves the Wikipedia page on Calculator, [^2] https://en.wikipedia.org/wiki/Calculator but it does not contain information about the AC button; (ii) Spelling mistakes in query generation. We construct the NoMIRACL non-relevant subset with queries with all top-$k$ passages judged as non-relevant by an annotator with a relevance of 00.
Relevant subset. Queries with at least one relevant passage are included in MIRACL Zhang et al. (2023). These queries contain known answers, as at least one of top-$k$ retrieved passages is marked relevant to provide sufficient information to answer the query. We construct the NoMIRACL relevant subset with queries with at least one relevant passage, i.e., all query-passage pairs have been explicitly judged by an annotator as either relevant with a relevance of $1$ or non-relevant with relevance of 00.
@@ 3.2 Fold Creation
3.3 No Available Answers
Coming up with reliable and accurate answers for a given query requires domain experts as annotators. Annotators can potentially highlight short extractive spans of answers within relevant passages, however, non-extractive queries can either contain multiple answers or a long-form answer, making it difficult to highlight a relevant answer span. Therefore, for NoMIRACL, we focus on the evaluation of top-$k$ passages as information contexts, which were easily judged for their relevancy by a data annotator.
@@ 3.4 Languages Covered and Dataset Usage
NoMIRACL covers 18 diverse typological languages Zhang et al. (2023). We cover 10 language families (from Niger-Congo to Indo European), 11 scripts (from Latin to Devanagari) covering diversity from the perspective of linguistic characteristics. The languages along with their ISO codes are listed below: Arabic (ar), Bengali (bn), German (de), English (en), Spanish (es), Persian (fa), Finnish (fi), French (fr), Hindi (hi), Indonesian (id), Japanese (ja), Korean (ko), Russian (ru), Swahili (sw), Thai (th), Yoruba (yo), Chinese (zh).
NoMIRACL is not a static dataset. From Table 1, we observe an uneven amount of queries present in each language. To avoid this, at runtime, a user provides a sampling ratio for queries in the non-relevant to relevant subset to utilize the maximum amount of queries available in NoMIRACL. Using the sampling ratio, we compare each subset (non-relevant and relevant) and return the maximum amount of queries possible.
@@ 4 Experimental Setup
Question-answering datasets, such as Natural Questions Kwiatkowski et al. (2019) or WebQuestions Berant et al. (2013) include a well-defined extractive answer for a majority of queries present within the dataset. In contrast, NoMIRACL does not contain answers as highlighted in (§3.3), but rather human-judged relevance judgements for all top-$k$ retrieved passages for 18 different languages. Therefore, we wish to evaluate LLM robustness, i.e., whether the query can be answered using the retrieved passages in RAG.
@@ 4.1 Evaluation Setup and Metrics
In NoMIRACL, we evaluate LLM hallucination using an input query, a vanilla prompting technique and top-$k$ (oracle) retrieved passages.
@@
Retrieved Passages. For each query in NoMIRACL, a maximum of $k=10$ passages are retrieved and judged by our annotators. We follow the hybrid retrieval setup in Zhang et al. (2023), which includes three different multilingual retriever models: (i) BM25 Robertson and Zaragoza (2009), a lexical retriever, previously shown to be robust across domains and languages Thakur et al. (2021); Zhang et al. (2022). We use the BM25 implementation available in Anserini Yang et al. (2018) with default parameters and the corresponding language-specific analyzer. (ii) mDPR Karpukhin et al. (2020), a dense retriever, using mBERT Devlin et al. (2018) as the backbone and fine-tuned on MS MARCO with the Tevatron toolkit Gao et al. (2023a). (iii) mColBERT Khattab and Zaharia (2020), a multi-vector retriever, fine-tuned following Khattab and Zaharia (2020) using mBERT as backbone and fine-tuned on MS MARCO. The top-$k$ passages are ranked using an ensemble fusion by normalizing each individual model score within the range $[0,1]$ and further averaging the scores.
@@
Robustness Evaluation. In our work, we evaluate LLM robustness as a response string $y$ in a binary classification task. Following prior evaluation strategies in Adlakha et al. (2023); Yu et al. (2023), we use the input query $q_{i}$, a vanilla prompt template $Q$, and a set of top-$k$ annotated passages $P_{k}$. We prompt the LLM to evaluate whether $q_{i}$ can be answered using any passage in $P_{k}$. The LLM generates a binary output containing either $y$ = “I don’t know” as $\mathrm{N}$ or “Yes, answer is present” as $\mathrm{P}$. The output is tagged “Invalid” if it does not fall in either one of the above. Recall from (§2.1), we calculate the hallucination rate (in %) = $\mathrm{FP}/\mathrm{FP}+\mathrm{FN}$, which measures the error in rejecting information from non-relevant passages and the error rate (in %) = $\mathrm{TN}/\mathrm{TN}+\mathrm{TP}$, which measures the error in identifying relevant passages amongst noisy ones.
4.2 Baseline Model: GPT-4
NoMIRACL provides a baseline based on a closed-book LLM: GPT-4 OpenAI (2023). GPT-4 has been trained to follow an instruction via a prompting strategy Ouyang et al. (2022) and provides a detailed response. Our choice lies with GPT-4 for its multilingual capabilities in machine translation Jiao et al. (2023), and natural language inference tasks Ahuja et al. (2023).

### 04 Mar 2024 v2

@@ Abstract
Retrieval-augmented generation (RAG) grounds large language model (LLM) output by leveraging external knowledge sources to reduce factual hallucinations. However, prior works lack a comprehensive evaluation of different language families, making it challenging to evaluate LLM robustness against errors in external retrieved knowledge. To overcome this, we establish NoMIRACL, [^1] NoMIRACL: https://huggingface.co/datasets/miracl/nomiracl a human-annotated dataset for evaluating LLM robustness in RAG across 18 typologically diverse languages. NoMIRACL includes both a non-relevant and a relevant subset. Queries in the non-relevant subset contain passages judged as non-relevant, whereas queries in the relevant subset include at least a single judged relevant passage. We measure LLM robustness using two metrics: (i) hallucination rate, measuring model tendency to hallucinate an answer, when the answer is not present in passages in the non-relevant subset, and (ii) error rate, measuring model inaccuracy to recognize relevant passages in the relevant subset. In our work, we measure robustness for a wide variety of multilingual-focused LLMs and observe that most of the models struggle to balance the two capacities. Models such as LLAMA-2, Orca-2, and FLAN-T5 observe more than an 88% hallucination rate on the non-relevant subset, whereas, Mistral overall hallucinates less, but can achieve up to a 74.9% error rate on the relevant subset. Overall, GPT-4 is observed to provide the best tradeoff on both subsets, highlighting future work necessary to improve LLM robustness. [^2] Baselines and model evaluation code are available at: https://github.com/project-miracl/nomiracl
@@ 1 Introduction
Retrieval Augmented Generation (RAG) (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021; Borgeaud et al., 2022) is a promising way to incorporate external knowledge via a first-stage retrieval system. RAG instills information from reliable knowledge corpora (provided as external passages) to generate accurate and faithful responses Shuster et al. (2021); Gao et al. (2023b); Liu et al. (2023). Ever since the advent of large language models (LLMs), such as GPT-3 Brown et al. (2020) or LLAMA-2 Touvron et al. (2023), they are the de-facto choice for generation in RAG, due to their unprecedented progress in text generation and understanding Brown et al. (2020); Li et al. (2021); Chang et al. (2023); Guo et al. (2023). RAG grounds the LLM-generated output, thereby avoiding previously observed factual hallucination Maynez et al. (2020); Raunak et al. (2021) and outdated knowledge Cao et al. (2021); He et al. (2023) in LLMs.
@@
Figure 1: LLM robustness evaluation as a binary tree in NoMIRACL. When dealing with queries in the non-relevant subset, the LLM is expected to disregard all noisy passages and refrain from answering ($\mathrm{TN}$). Conversely, for queries in the relevant subset, the LLM should recognize the relevant passage and provide a valid answer ($\mathrm{TP}$).
@@
A challenging issue in RAG is to provide robust and reliable LLM-generated output. The generation stage is dependent on the first-stage information retrieval system. The retrieval system poses challenges in accurately retrieving relevant information when evaluated either on zero-shot domains Thakur et al. (2021) or low-resource languages Zhang et al. (2023). The incorrect or non-relevant information contained in retrieved passages can frequently mislead the LLM to hallucinate Adlakha et al. (2023); Chen et al. (2023); Shi et al. (2023); Yoran et al. (2023); Yu et al. (2023). Prior work lacks a comprehensive evaluation of LLM reasoning capabilities in multiple languages. As a result, it still remains unclear to which extent LLMs hallucinate across both high- or low-resource languages.
@@
To facilitate research in this direction, we present NoMIRACL, a large multilingual human-annotated dataset containing over 56,000 (including both non-relevant and relevant samples) to evaluate LLM robustness against errors in first-stage external information, i.e., retrieved passages, across 18 typologically diverse languages. To construct the dataset, we hired 31 native speakers as human annotators Zhang et al. (2023). NoMIRACL contains two subsets, non-relevant and relevant. The non-relevant subset contains all queries with no-known answers, i.e., all top-$k$ retrieved passages manually judged as non-relevant. Conversely, the relevant subset contains queries with known answers, i.e., at least one of the top-$k$ passages is manually judged as relevant.
To better understand the LLM robustness in NoMIRACL, we conduct experiments with several existing powerful and multilingual-focused LLMs (e.g., GPT-4, Mistral, LLAMA-2). We conduct our experiments using the top-$k$ oracle passages retrieved using a hybrid retrieval system from a language-specific Wikipedia corpus Zhang et al. (2023). We use a zero-shot “vanilla” prompt template for prompting all LLMs. Our key findings are as follows: First, a majority of LLMs such as LLAMA-2, Orca-2, and FLAN-T5 observe a suprisingly high 88% hallucination rate on the non-relevant subset. Second, Mistral deals with hallucination well but performs poorly on the relevant subset. Overall, GPT-4 is found to provide the optimal performance tradeoff across both subsets.
Additionally, we conduct an empirical analysis on the complete NoMIRACL English to analyze the blind spots of LLMs, where we observe LLAMA-2-7B and 13B are interestingly found to repeat the query and prompt instructions on average by at least 25%. Mistral-7B always provides a rationale in their output generation by over 88%. This empirical analysis can be used to discover hallucination and error patterns and develop effective methods to alleviate them in LLMs.
@@
Figure 2: Confusion matrix for robustness evaluation with NoMIRACL. More details are provided in (§2.1); (Subset) denotes the ground-truth in NoMIRACL; (Pred.) denotes the LLM output prediction.
@@
To summarize, our contributions are: (i) We introduce NoMIRACL, a novel multilingual dataset to evaluate LLM hallucinations against first-stage retrieval errors in RAG. (ii) We evaluate several powerful multilingual LLMs and observe challenges in LLM robustness by often hallucinating an answer within non-answerable passages in the non-relevant subset and the inability to recognize relevant passages in the relevant subset. (iii) We conduct thorough manual inspections on LLM’s generation results, and find several hallucination patterns for each genre of the LLM; We hope NoMIRACL can serve as a valuable dataset towards a much-needed LLM robustness evaluation.
@@ 2 Background and Problem Identification
A challenging issue in RAG is to provide robust and reliable LLM-generated output against a first-stage information retrieval system. Overreliance of the LLM on the content of the retrieved passages (i.e., tendency to extract information from passages) can be limiting when passages are noisy or non-relevant Yu et al. (2023); Yoran et al. (2023); Shi et al. (2023).
@@ 2.1 Robustness Evaluation
We conduct our evaluation strategy as a contingency table (as shown in Figure 2) to robustly evaluate the LLM behavior in both answerable and non-answerable scenarios using a binary classification task, by comparing LLM predictions against the ground truth provided by human annotators.
@@
Confusion Matrix. In our confusion matrix (cf. Figure 2), for our non-relevant subset, True Negative ($\mathrm{TN}$) denotes when the model correctly predicts queries with no-known answers using the non-relevant retrieved passages, whereas False Positive ($\mathrm{FP}$) denotes when the model prediction is incorrect on the non-relevant subset. Similarly, True Positive ($\mathrm{TP}$) denotes when LLM correctly predicts queries with known answers within the retrieved passages, whereas False Negative ($\mathrm{FN}$) denotes when the model prediction is incorrect (i.e., the model finds no answer) on the relevant subset.
Evaluation metrics. Following prior works Adlakha et al. (2023); Chen et al. (2023), we assess LLM robustness as a binary classification task using two metrics: (i) hallucination rate and (ii) error rate. First, we compute the hallucination rate (in %) $=\mathrm{FP}/(\mathrm{FP}+\mathrm{TN})$ which measures the model’s tendency to hallucinate an answer, when no answer is available in all of the passages in the non-relevant subset. Next, we measure the error rate (in %) $=\mathrm{FN}/(\mathrm{FN}+\mathrm{TP})$ which measures the model’s ability to identify the answer present within the passages in the relevant subset.
@@ 3 NoMIRACL Dataset
As the goal of NoMIRACL is to understand to which extent LLMs tend to hallucinate across different languages, our dataset contains 18 diverse languages with a myriad of both correct or answerable queries (relevant) subset and hallucinated or unanswerable queries (non-relevant). We describe our dataset construction procedure in (§3.1), fold creation in (§3.2) and languages covered and dataset usage in (§3.3). An overview of our data construction procedure is shown in Figure 3.
@@ 3.1 Data Construction Procedure
NoMIRACL, a supplementary dataset, was created in the course of developing MIRACL Zhang et al. (2023). NoMIRACL data construction occured in two stages, following Zhang et al. (2023); Clark et al. (2020). In the first stage, the annotator (a native language speaker) is asked to generate a well-formed query for each individual prompt text. Each prompt is a short text snippet containing the first 100 words from a language-specific Wikipedia corpus. Next, for each human-generated query, top-$k$ passages are retrieved from the Wikipedia corpus using a hybrid multilingual retrieval system (more details in §4.1). In the second stage, annotators assess the relevance of the top-$k$ query–passage pairs as a binary judgment, either relevant (relevance = $1$) or non-relevant (relevance = 00). For additional details in data construction, such as quality control, we would like the refer the reader to Appendix C.
Non-relevant subset. Annotators generate queries based on certain dataset guidelines, however, occasionally the human-generated queries cannot be answered with the external corpus, which leads to the scenario where none of the top-$k$ passages are relevant, i.e., none contains the answer. These queries with unknown answers may occur due to the following reasons: (i) queries can be either generic or specific for information to be present in Wikipedia, for e.g. “What does the AC button on a calculator stand for?” retrieves the Wikipedia page on Calculator, [^3] https://en.wikipedia.org/wiki/Calculator but it does not contain information about the AC button; (ii) spelling mistakes in query generation. We construct the non-relevant subset with queries with all top-$k$ passages judged as non-relevant, i.e., with a relevance of 00.
Relevant subset. Queries with known answers, i.e. at least one of top-$k$ retrieved passages is marked by the annotator as relevant to provide sufficient information to answer the query. We construct the NoMIRACL relevant subset with queries with at least one relevant passage, i.e., all query-passage pairs have been judged as either relevant with a relevance of $1$ or non-relevant with relevance of 00.
@@ 3.3 Languages Covered and Dataset Usage
NoMIRACL covers 18 diverse typological languages Zhang et al. (2023). We cover 10 language families (from Niger-Congo to Indo-European), and 11 scripts (from Latin to Devanagari) covering diversity from the perspective of linguistic characteristics. The languages along with their ISO codes are listed below: Arabic (ar), Bengali (bn), German (de), English (en), Spanish (es), Persian (fa), Finnish (fi), French (fr), Hindi (hi), Indonesian (id), Japanese (ja), Korean (ko), Russian (ru), Swahili (sw), Thai (th), Yoruba (yo), Chinese (zh).
From Table 1, we observe an uneven amount of queries present in each language. To avoid this non-uniformity and due to additional budget constraints (see subsection 4.3), in our experiments, we limit the maximum number of 250 queries for each language (if available) and subset in NoMIRACL.
@@ 4.1 Evaluation Setup and Metrics
In NoMIRACL, we evaluate LLM hallucination or error, using an input query, a vanilla prompting technique and top-$k$ (oracle) retrieved passages.
@@
Retrieved Passages. For each query in NoMIRACL, a maximum of $k=10$ passages are retrieved and judged by our annotators. We follow the hybrid retrieval setup in Zhang et al. (2023), which includes three different multilingual retriever models: (i) BM25 Robertson and Zaragoza (2009), a lexical retriever, previously shown to be robust across domains and languages Thakur et al. (2021); Zhang et al. (2022). We use the BM25 implementation available in Anserini Yang et al. (2018) with default parameters ($k_{1}=0.9$ and $b=0.4$) and the corresponding language-specific analyzer. (ii) mDPR Karpukhin et al. (2020), a dense retriever, using mBERT Devlin et al. (2018) as the backbone and fine-tuned on MS MARCO with the Tevatron toolkit Gao et al. (2023a). (iii) mColBERT Khattab and Zaharia (2020), a multi-vector retriever, fine-tuned following Khattab and Zaharia (2020) using mBERT as backbone and fine-tuned on MS MARCO. The top-$k$ passages are ranked using an ensemble fusion by normalizing and averaging each model score within the range of $[0,1]$.
@@
Evaluation Objective. In our work, we evaluate LLM robustness as a response string $y$ in a binary classification task. Following prior evaluation strategies in Adlakha et al. (2023); Yu et al. (2023), we use the input query $q_{i}$, a vanilla prompt template $Q$, and a set of top-$k$ annotated passages $P_{k}$. We prompt the LLM to evaluate whether $q_{i}$ can be answered using any passage in $P_{k}$. The LLM generates a binary output containing either $y$ = “I don’t know” as $\mathrm{N}$ or “Yes, answer is present” as $\mathrm{P}$. The output is tagged “Invalid” if it does not fall in either one of the above. Recall from (§2.1), we calculate the hallucination rate (in %) = $\mathrm{FP}/(\mathrm{FP}+\mathrm{TN})$, which measures the error in rejecting information from non-relevant passages and the error rate (in %) = $\mathrm{FN}/(\mathrm{FN}+\mathrm{TP}$), which measures the error in identifying relevant passages amongst noisy ones.
4.2 Evaluation Models
We evaluate eleven state-of-the-art LLMs with a strong focus of multilingual instruction capabilities, including both open and closed-sourced. All model checkpoints can be found in Table 3.
(1) OpenAI: We include both closed-book LLM variants: GPT-3.5-turbo and GPT-4 OpenAI (2023) using the Azure OpenAI service. (2) Mistral: We include two variants: (i) Mistral-7B-Instruct-v0.2, a 7B instruction-tuned parameter model Jiang et al. (2023) and (ii) Mixtral-8x7B-v0.1, a sparse ixture-of-Expert (MoE) model Jiang et al. (2024). (3) Orca-2: In the Orca-2 series Mitra et al. (2023), we include both Orca-2-7B and Orca-2-13B. (4) Aya: Aya-101 Üstün et al. (2024) is a recently introduced multilingual LLM containing 13B parameters and trained with 101 languages. (5) LLAMA-2: In the LLAMA-2 series Touvron et al. (2023), we include three chat variants: Llama-2-7b-chat-hf, Llama-2-13b-chat-hf and Llama-2-70b-chat-hf instruction tuned chat models. (6) FLAN-T5: One of the earliest instruction-tuned LMs, we include flan-T5-xxl containing 11B parameters Chung et al. (2022).
4.3 Experimental Settings